# LUMINA: A Multi-Vendor Mammography Benchmark with Energy Harmonization Protocol

### 15 Mar 2026 v1

@@ Abstract
Publicly available full-field digital mammography (FFDM) datasets remain limited in size, clinical labels, and vendor diversity, which hinders the training of robust models. We present LUMINA, a curated, multi-vendor FFDM dataset that explicitly encodes acquisition energy and vendor metadata to expose clinically relevant appearance shifts that current benchmarks overlook. This innovative resource comprises 1824 images from 468 patients (960 benign, 864 malignant) with pathology-confirmed outcomes, BI-RADS assessments, and breast-density annotations. LUMINA spans six acquisition systems and both high- and low-energy styles, exposing vendor- and energy-driven appearance shifts. To reduce cross‑vendor/energy drift while preserving lesion morphology, we introduce a foreground‑only, pixel‑space alignment (“energy harmonization”) that aligns each image to a low‑energy reference style, leaving the zero‑valued background unchanged. By benchmarking modern CNN and transformer baselines on three clinically meaningful tasks—diagnosis (benign vs. malignant), BI‑RADS risk grouping, and density—we unify single‑vs‑two‑view evaluation and show that two‑view models consistently outperform single‑view; in our benchmark, EfficientNet‑B0 attains AUC 93.54% for diagnosis, and Swin‑T yields the best macro‑AUC 89.43% for density. Harmonization improves AUC/ACC across backbones and yields more focal Grad‑CAM localization around suspicious regions. Being a richly annotated resource, LUMINA thus provides (a) a vendor‑diverse, energy‑labeled benchmark and (b) a model‑agnostic harmonization protocol that together catalyze reliable, deployable mammography AI.
@@ 1 Introduction
Breast cancer is the most prevalent type of cancer and one of the leading causes of cancer-related deaths among women [18, 34]. Recent advances in deep learning have demonstrated the potential to improve diagnostic accuracy and detect subtle lesions that can be overlooked by human readers [30, 17]. Training robust and generalizable models, however, requires large-scale, high-quality datasets. In the mammography domain, publicly available datasets such as CBIS-DDSM [14] and INbreast [20] are limited in size and heterogeneity, restricting the development of artificial intelligence (AI) systems that can perform reliably across diverse clinical settings. This highlights the critical need for comprehensive mammography datasets that capture high-resolution images across multiple views, patient populations, and imaging systems. To this end, we introduce the LUMINA dataset and a foreground-only pixel-space CDF alignment that reduces vendor/energy appearance drift and consistently improves accuracy and AUC across three tasks.
@@
Foreground-only histogram harmonization. A simple, vendor-agnostic, object-masked histogram matching that excludes background (intensity 0), aligns to a low-energy reference, and preserves lesion contrast; formalized with CDF mapping.
@@ 4 Pixel-Space Harmonization
Modern FFDM systems typically acquire images using either low-energy or high-energy X-ray settings. Low-energy mammograms are optimized for soft-tissue contrast, enhancing the visibility of subtle lesions, while high-energy mammograms emphasize denser structures such as calcification. However, these differing acquisition styles, combined with variations in vendor-specific post-processing algorithms, result in substantial intensity and contrast discrepancies across devices. Such inconsistencies can degrade the generalization performance of deep learning models, as models trained on one vendor’s imaging style may perform poorly on another. Therefore, harmonizing intensity distributions across imaging systems is critical for developing vendor-agnostic and clinically reliable AI models.
@@
where the sums start from intensity 1 because intensity 0 corresponds to background and is excluded. The mapping function is defined by matching these CDFs:
@@ 6.2 Benchmark Results
Breast cancer diagnosis: Table 4 shows that EfficientNet-B0 consistently achieved higher AUC than ResNet-50 and DenseNet-121, demonstrating its strong capability for mammographic image analysis. For instance, single-view EfficientNet-B0 achieves 90.66% AUC at and 92.13% AUC at , outperforming both ResNet-50 (87.42% / 86.69%) and DenseNet-121 (87.14% / 90.52%). Higher input resolutions generally improve performance, reflecting the importance of preserving fine mammography details. Nevertheless, inputs still provide competitive results, offering computational efficiency. For example, single-view EfficientNet-B0 has 4.01M parameters and 0.41G FLOPs at , compared to 4.01M parameters and 2.13G FLOPs at . Two-view models outperform single-view models across all backbones, highlighting the value of combining CC and MLO views. Notably, two-view EfficientNet-B0 with inputs achieves the highest overall performance, reaching an AUC of 93.54%, representing the most effective configuration in our benchmark. Fig. 5 illustrates the t-SNE visualization, which shows how EfficientNet-B0 learned representations separate benign and malignant cases more clearly at higher input resolution.
@@ 7 Conclusion
We present LUMINA, a curated dataset of full-field digital mammograms with comprehensive pathology-confirmed, BI-RADS, and breast density annotations. We further proposed a foreground-only histogram method to mitigate inter-scanner intensity variations, improving consistency across imaging systems. Using this harmonized dataset, we systematically benchmarked several state-of-the-art deep learning models on three clinically meaningful tasks: breast cancer diagnosis, BI-RADS classification, and breast density assessment. Experimental results demonstrated consistent performance gains after harmonization, underscoring the importance of vendor-level normalization for robust mammography AI. We expect LUMINA to create a practical bridge between academic benchmarks and deployment‑context generalization for breast cancer.
@@ 8 Data and Code
The LUMINA dataset is available at https://osf.io/b63jc/, and our code is available at https://github.com/NUBagciLab/LUMINA.
9 Acknolodgement
This research was funded by NIH grant number R01-HL171376.
@@ 9 Acknolodgement
Table 9 and summarizes the unified hyperparameter settings and pre-processing operations used across all model architectures, respectively. These settings and operations ensured a controlled and fair comparison in our benchmark experiments.

### 17 Mar 2026 v2

@@ Abstract
Publicly available full-field digital mammography (FFDM) datasets remain limited in size, clinical annotations, and vendor diversity, hindering the development of robust models. We introduce LUMINA, a curated, multi-vendor FFDM dataset that explicitly encodes acquisition energy and vendor metadata to capture clinically relevant appearance variations often overlooked in existing benchmarks. This dataset contains 1824 images from 468 patients (960 benign, 864 malignant), with pathology-confirmed labels, BI-RADS assessments, and breast-density annotations. LUMINA spans six acquisition systems and includes both high- and low-energy imaging styles, enabling systematic analysis of vendor- and energy-induced domain shifts. To address these variations, we propose a foreground-only pixel-space alignment method (“energy harmonization”) that maps images to a low-energy reference while preserving lesion morphology. We benchmark CNN and transformer models on three clinically relevant tasks: diagnosis (benign vs. malignant), BI-RADS classification, and density estimation. Two-view models consistently outperform single-view models. EfficientNet-B0 achieves an AUC of 93.54% for diagnosis, while Swin-T achieves the best macro-AUC of 89.43% for density prediction. Harmonization improves performance across architectures and produces more localized Grad-CAM responses. Overall, LUMINA provides (1) a vendor-diverse benchmark and (2) a model-agnostic harmonization framework for reliable and deployable mammography AI.
@@ 1 Introduction
Breast cancer is the most prevalent type of cancer and one of the leading causes of cancer-related deaths among women [18, 34]. Recent advances in deep learning have demonstrated the potential to improve diagnostic accuracy and detect subtle lesions that can be overlooked by human readers [30, 17]. However, training robust and generalizable models requires large-scale, high-quality datasets. In the mammography domain, publicly available datasets such as CBIS-DDSM [14] and INbreast [20] are limited in size and heterogeneity, restricting the development of artificial intelligence (AI) systems that can perform reliably across diverse clinical settings. This underscores the need for comprehensive mammography datasets that capture high-resolution images across multiple views, patient populations, and imaging systems. To address this gap, we introduce the LUMINA dataset along with a foreground-only pixel-space CDF alignment method that reduces vendor/energy appearance drift and consistently improves accuracy and AUC across three tasks.
@@
Foreground-only histogram harmonization. A simple, vendor-agnostic histogram matching that excludes background pixels and preserves lesion contrast.
@@ 4 Pixel-Space Harmonization
Modern FFDM systems typically acquire images using either low-energy or high-energy X-ray settings. Low-energy mammograms are optimized for soft-tissue contrast, enhancing the visibility of subtle lesions, while high-energy mammograms emphasize denser structures such as calcification. However, these acquisition differences, combined with vendor-specific processing, introduce significant intensity and contrast variations across devices. Such inconsistencies can degrade the generalization performance of deep learning models, as models trained on one vendor’s imaging style may perform poorly on another. Therefore, harmonizing intensity distributions across systems is critical for developing vendor-agnostic and clinically reliable AI models.
@@
where the summation starts from intensity 1, as intensity 0 represents the background and is excluded. The mapping function is defined by matching these CDFs:
@@ 6.2 Benchmark Results
Breast cancer diagnosis: Table 4 shows that EfficientNet-B0 consistently outperformed ResNet-50 and DenseNet-121 in terms of AUC, indicating its strong capability for mammographic image analysis. For instance, single-view EfficientNet-B0 achieves 90.66% AUC at and 92.13% AUC at , outperforming both ResNet-50 (87.42% / 86.69%) and DenseNet-121 (87.14% / 90.52%). Higher input resolutions generally improve performance, highlighting the importance of preserving fine mammography details. Nevertheless, inputs still provide competitive results, offering computational efficiency. For example, single-view EfficientNet-B0 has 4.01M parameters and 0.41G FLOPs at , compared to 4.01M parameters and 2.13G FLOPs at . Two-view models outperform single-view models across all backbones, highlighting the value of combining CC and MLO views. Notably, two-view EfficientNet-B0 with inputs achieves the highest overall performance, reaching an AUC of 93.54%, representing the most effective configuration in our benchmark. Fig. 5 illustrates the t-SNE visualization, which shows how EfficientNet-B0 learned representations separate benign and malignant cases more clearly at higher input resolution.
@@ 7 Conclusion
We present LUMINA, a multi-vendor FFDM dataset with comprehensive pathology, BI-RADS, and density annotations. We also introduced a foreground-only histogram-based harmonization method to mitigate inter-scanner variability. Extensive experiments demonstrate that harmonization consistently improves performance across multiple tasks and architectures. These findings highlight the importance of vendor-level normalization for robust mammography AI. We expect LUMINA to bridge the gap between academic benchmarks and real-world clinical deployment.
@@ 8 Data and Code
The LUMINA dataset and the associated source code are publicly available at the following links:
OSF (Dataset): https://osf.io/b63jc/
Kaggle (Dataset): https://www.kaggle.com/datasets/phy710/lumina-mammography-dataset
GitHub (Code): https://github.com/NUBagciLab/LUMINA
9 Acknowledgment
This research was partially funded by NIH: R01-HL171376.
@@ 9 Acknowledgment
Table 9 summarizes the unified hyperparameter settings and pre-processing operations used across all model architectures, respectively. These settings and operations ensured a controlled and fair comparison in our benchmark experiments.